# Chinese Academy of Sciences, Beijing 100190, P. R. China

### v1

@@ Appendix B Surface energy-momentum tensor
In the wall rest frame $K^{\prime}$, the surface tensor is given by Eqs. (75) and (76). In the observer’s rest frame $K$, $K^{\prime}$ moves along the $\mathbf{x}$ axis with velocity $\overline{v}_{w}$, then the surface density tensor $S_{\mu\nu}^{x}(\overline{v}_{w})$ measured by the observer is obtained by performing the Lorentz transformation
$S_{00}^{x}$ $=\gamma^{2}(S_{00}^{\prime}+2\overline{v}_{w}S_{01}^{\prime}+{\overline{v}_{w}}^{2}S_{11}^{\prime}),$ (77)
@@
$S_{\mu\nu}^{x}=\left(\begin{array}[]{cccc}\gamma^{2}\sigma_{w}&\gamma^{2}\sigma_{w}\overline{v}_{w}&0&0\\ \gamma^{2}\sigma_{w}\overline{v}_{w}&\gamma^{2}\sigma_{w}\overline{v}_{w}^{2}&0&0\\ 0&0&-\sigma_{w}&0\\ 0&0&0&-\sigma_{w}\end{array}\right),$ (78)
$S_{\mu\nu}^{y}=\left(\begin{array}[]{cccc}\gamma^{2}\sigma_{w}&0&\gamma^{2}\sigma_{w}\overline{v}_{w}&0\\ 0&-\sigma_{w}&0&0\\ \gamma^{2}\sigma_{w}\overline{v}_{w}&0&\gamma^{2}\sigma_{w}\overline{v}_{w}^{2}&0\\ 0&0&0&-\sigma_{w}\end{array}\right),$ (79)
$S_{\mu\nu}^{z}=\left(\begin{array}[]{cccc}\gamma^{2}\sigma_{w}&0&0&\gamma^{2}\sigma_{w}\overline{v}_{w}\\ 0&-\sigma_{w}&0&0\\ 0&0&-\sigma_{w}&0\\ \gamma^{2}\sigma_{w}\overline{v}_{w}&0&0&\gamma^{2}\sigma_{w}\overline{v}_{w}^{2}\end{array}\right).$ (80)
@@
$S_{\mu\nu}(\overline{v}_{w})=\frac{\sigma_{w}}{3}{\rm diag}(\gamma^{2},\overline{v}_{w}^{2}\gamma^{2}-2,\overline{v}_{w}^{2}\gamma^{2}-2,\overline{v}_{w}^{2}\gamma^{2}-2).$ (81)
@@ Appendix C Wall evolution model
We follow Ref. Martins:2016book to review the VOS model for describing the evolution of the wall. Let us start from the world volume (Dirac) action Martins:2016book
$S=-\int\mathcal{L}d^{3}\sigma=-\sigma_{w}\int\sqrt{\gamma}d^{3}\sigma,$ (82)
@@
$\frac{\dot{a}}{a}\delta_{0\lambda}\sqrt{\gamma}\gamma^{ab}\gamma_{ab}-\partial_{c}\left(\sqrt{\gamma}\gamma^{ab}g_{\mu\lambda}x_{,a}^{\mu}\delta_{b}^{c}\right)=0.$ (83)
@@
$\bar{\rho}_{w}\equiv\frac{E}{V}=\frac{\sigma_{w}a^{2}}{V}\int\varepsilon d^{2}\sigma,\quad\overline{v}_{w}^{2}\simeq v_{\rm rms}^{2}=\frac{\int\dot{x}^{2}\varepsilon d^{2}\sigma}{\int\varepsilon d^{2}\sigma},$ (84)
where $\varepsilon=\sqrt{\gamma}\gamma^{00}/a$ with $a$ being the cosmic scale factor. By averaging the temporal and spatial components of the EOM (83), we obtan Martins:2016book
$\frac{d\overline{\rho}_{w}}{dt}$ $=-H\overline{\rho}_{w}\left(1+3\overline{v}_{w}^{2}\right)=-3(1+w)H\overline{\rho}_{w},$ (85)
@@
where $t$ is the physical time, $H=\dot{a}/a$ is the Hubble parameter, and the momentum parameter $k_{w}=0.66\pm 0.04$ from the simulations Martins:2016ois . It is assumed in Eq. (85) that the curvature radii are identical to the correlation length $\overline{L}$. Taking the one-scale assumption $\overline{\rho}_{w}=\sigma_{w}/\overline{L}$ and adding the energy loss processes (damping effects) by hand to Eq. (85), one obtains the VOS model for the description of domain wall as Martins:2016book
$\frac{d\overline{L}}{dt}$ $=H\overline{L}+\frac{\overline{L}}{\ell_{d}}\overline{v}_{w}^{2}+c_{w}\overline{v}_{w},$ (86)
@@
where the chopping parameter $c_{w}=0.81\pm 0.04$ from the simulations of domain wall evolution in the $Z_{2}$ model Martins:2016ois . The damping length is defined as
$\frac{1}{\ell_{d}}=3H+\frac{1}{\ell_{f}},$ (87)
which includes the damping effects on the network from Hubble drag and particle friction, which is related to the particle pressure $\Delta P$ by
$\frac{1}{\ell_{f}}=\frac{\Delta P}{\overline{v}_{w}\sigma_{w}}.$ (88)
@@ Appendix D Beta function
$\beta(x)=\gamma_{p}{\rm erfc}(x),$ (89)
@@
We plot $\beta(x)$ as a function of $x=\delta_{c}/(\sqrt{2}\sigma)$ in Fig. 10. We observe that to obtain an acceptable PBH abundance, $x$ should fall in the range of $3-5$ to exponentially suppress the PBH production. Therefore, the formation of PBHs via the critical density collapse mechanism arises from the tail of the probability density function.
@@ Appendix E Strategy for PTA
In this section, we provide further details of the PTA data analysis. We follow the analysis procedure encoded in PTArcade, which wraps enterprise and Ceffyl to allow us to perform a standard search strategy for the PTA datasets (see Ref. Mitridate:2023oar for more details). The searches for PTA stochastic signals adopt the pulsars’ timing residuals, which are modeled as
$\overrightarrow{\delta t}=\vec{n}+\bm{F}\vec{a}+\bm{M}\vec{\epsilon},$ (92)
where the terms on the right-hand side of Eq. (92) are the white noise, red noise, and small errors in the fit to the timing-ephemeris parameter NANOGrav:2020tig , respectively.
Three types of white noise per backend/receiver system are included in the analysis, i.e., EFAC $E_{k}$, EQUAD $Q_{k}$[s], and ECORR $J_{k}$[s]. The pulsar-intrinsic red noise and GW background signals are included in the second term on the right-hand side of Eq. (92), which describes time-correlated stochastic processes. We adopt a power-law red noise with two parameters per pulsar for each dataset: the amplitude $A_{\rm red}$ and the spectral index $\gamma_{\rm red}$. For the GW signal, we adopt the domain wall annihilation scenario with four parameters in both datasets: the domain wall energy fraction $f_{w}(T_{\rm ann})$ at the annihilation temperature $T_{\rm ann}$, the domain wall annihilation temperature $T_{\rm ann}$ in units of GeV, the spectral index above the peak frequency $c_{1}$, and the parameter that represents the width of the spectral index $c_{2}$. The spectral $\gamma_{c}$ is fixed at 3 to respect the causality. In table 3, we provide the priors for the parameters of white and red noises, as well as the priors for parameters of the GW background signal from the annihilation of the domain walls.
@@
[…]
$Q_{k}$[s] EQUAD per backend/receiver system Log-uniform $[-8.5,\leavevmode\nobreak\ -5]$ Single-pulsar analysis only
$J_{k}$[s] ECORR per backend/receiver system Log-uniform $[-8.5,\leavevmode\nobreak\ -5]$ Single-pulsar analysis only
@@
$A_{\rm red}$ Red noise power-law amplitude Log-uniform $[-20,\leavevmode\nobreak\ -11]$ one parameter per pulsar
[…]
@@
$f_{w}(T_{\rm ann})$ Domain wall energy fraction at $T_{\rm ann}$ Log-Uniform $[-2,0]$ one parameter per PTA
@@
Table 4: Bayes estimators for the fit to NG15 and IPTA-DR2 datasets.
NG15 IPTA-DR2
${\rm Log}_{10}f_{w}(T_{\rm ann})$ $-0.71\pm 0.17$ $-0.93\pm 0.10$
${\rm Log}_{10}[T_{\rm ann}/\rm GeV]$ $-0.70\pm 0.19$ $-1.28\pm 0.15$
$c_{1}$ 0.95$\pm$0.29 0.97$\pm$0.29
$c_{2}$ 1.66$\pm$0.55 1.48$\pm$0.58

### v2

@@ Appendix B Surface energy-momentum tensor
In the wall rest frame $K^{\prime}$, the surface tensor is given by Eqs. (A.8) and (A.9). In the observer’s rest frame $K$, $K^{\prime}$ moves along the $\mathbf{x}$ axis with velocity $\overline{v}_{w}$, then the surface density tensor $S_{\mu\nu}^{x}(\overline{v}_{w})$ measured by the observer is obtained by performing the Lorentz transformation
$S_{00}^{x}$ $=\gamma^{2}(S_{00}^{\prime}+2\overline{v}_{w}S_{01}^{\prime}+{\overline{v}_{w}}^{2}S_{11}^{\prime}),$ (B.1)
@@
$S_{\mu\nu}^{x}=\left(\begin{array}[]{cccc}\gamma^{2}\sigma_{w}&\gamma^{2}\sigma_{w}\overline{v}_{w}&0&0\\ \gamma^{2}\sigma_{w}\overline{v}_{w}&\gamma^{2}\sigma_{w}\overline{v}_{w}^{2}&0&0\\ 0&0&-\sigma_{w}&0\\ 0&0&0&-\sigma_{w}\end{array}\right),$ (B.2)
$S_{\mu\nu}^{y}=\left(\begin{array}[]{cccc}\gamma^{2}\sigma_{w}&0&\gamma^{2}\sigma_{w}\overline{v}_{w}&0\\ 0&-\sigma_{w}&0&0\\ \gamma^{2}\sigma_{w}\overline{v}_{w}&0&\gamma^{2}\sigma_{w}\overline{v}_{w}^{2}&0\\ 0&0&0&-\sigma_{w}\end{array}\right),$ (B.3)
$S_{\mu\nu}^{z}=\left(\begin{array}[]{cccc}\gamma^{2}\sigma_{w}&0&0&\gamma^{2}\sigma_{w}\overline{v}_{w}\\ 0&-\sigma_{w}&0&0\\ 0&0&-\sigma_{w}&0\\ \gamma^{2}\sigma_{w}\overline{v}_{w}&0&0&\gamma^{2}\sigma_{w}\overline{v}_{w}^{2}\end{array}\right).$ (B.4)
@@
$S_{\mu\nu}(\overline{v}_{w})=\frac{\sigma_{w}}{3}{\rm diag}(3\gamma^{2},\overline{v}_{w}^{2}\gamma^{2}-2,\overline{v}_{w}^{2}\gamma^{2}-2,\overline{v}_{w}^{2}\gamma^{2}-2).$ (B.5)
@@ Appendix C Wall evolution model
We follow Ref. [47] to review the VOS model for describing the evolution of the wall. Let us start from the world volume (Dirac) action [47]
$S=-\int\mathcal{L}d^{3}\sigma=-\sigma_{w}\int\sqrt{\gamma}d^{3}\sigma,$ (C.1)
@@
$\frac{\dot{a}}{a}\delta_{0\lambda}\sqrt{\gamma}\gamma^{ab}\gamma_{ab}-\partial_{c}\left(\sqrt{\gamma}\gamma^{ab}g_{\mu\lambda}x_{,a}^{\mu}\delta_{b}^{c}\right)=0.$ (C.2)
@@
$\bar{\rho}_{w}\equiv\frac{E}{V}=\frac{\sigma_{w}a^{2}}{V}\int\varepsilon d^{2}\sigma,\quad\overline{v}_{w}^{2}\simeq v_{\rm rms}^{2}=\frac{\int\dot{x}^{2}\varepsilon d^{2}\sigma}{\int\varepsilon d^{2}\sigma},$ (C.3)
where $\varepsilon=\sqrt{\gamma}\gamma^{00}/a$ with $a$ being the cosmic scale factor. By averaging the temporal and spatial components of the EOM (C.2), we obtan [47]
$\frac{d\overline{\rho}_{w}}{dt}$ $=-H\overline{\rho}_{w}\left(1+3\overline{v}_{w}^{2}\right)=-3(1+w)H\overline{\rho}_{w},$ (C.4)
@@
where $t$ is the physical time, $H=\dot{a}/a$ is the Hubble parameter, and the momentum parameter $k_{w}=0.66\pm 0.04$ from the simulations [48]. It is assumed in Eq. (C.4) that the curvature radii are identical to the correlation length $\overline{L}$. Taking the one-scale assumption $\overline{\rho}_{w}=\sigma_{w}/\overline{L}$ and adding the energy loss processes (damping effects) by hand to Eq. (C.4), one obtains the VOS model for the description of domain wall as [47]
$\frac{d\overline{L}}{dt}$ $=H\overline{L}+\frac{\overline{L}}{\ell_{d}}\overline{v}_{w}^{2}+c_{w}\overline{v}_{w},$ (C.5)
@@
where the chopping parameter $c_{w}=0.81\pm 0.04$ from the simulations of domain wall evolution in the $Z_{2}$ model [48]. The damping length is defined as
$\frac{1}{\ell_{d}}=3H+\frac{1}{\ell_{f}},$ (C.6)
which includes the damping effects on the network from Hubble drag and particle friction, which is related to the particle pressure $\Delta P$ by [114]
$\frac{1}{\ell_{f}}=\frac{\Delta P}{\overline{v}_{w}\sigma_{w}}.$ (C.7)
@@ Appendix D Beta function
$\beta(x)=\gamma_{p}{\rm erfc}(x),$ (D.1)
@@
We plot $\beta(x)$ as a function of $x=\delta_{c}/(\sqrt{2}\sigma)$ in Fig. 11. We observe that to obtain an acceptable PBH abundance, $x$ should fall in the range of 353-5 to exponentially suppress the PBH production. Therefore, the formation of PBHs via the critical density collapse mechanism arises from the tail of the probability density function.
@@ Appendix E Strategy for PTA
In this section, we provide further details of the PTA data analysis. We follow the analysis procedure encoded in PTArcade, which wraps enterprise and Ceffyl to allow us to perform a standard search strategy for the PTA datasets (see Ref. [93] for more details). The searches for PTA stochastic signals adopt the pulsars’ timing residuals, which are modeled as
$\overrightarrow{\delta t}=\vec{n}+\bm{F}\vec{a}+\bm{M}\vec{\epsilon},$ (E.1)
where the terms on the right-hand side of Eq. (E.1) are the white noise, red noise, and small errors in the fit to the timing-ephemeris parameter [115], respectively.
Three types of white noise per backend/receiver system are included in the analysis, i.e., EFAC $E_{k}$, EQUAD $Q_{k}$[s], and ECORR $J_{k}$[s]. The pulsar-intrinsic red noise and GW background signals are included in the second term on the right-hand side of Eq. (E.1), which describes time-correlated stochastic processes. We adopt a power-law red noise with two parameters per pulsar for each dataset: the amplitude $A_{\rm red}$ and the spectral index $\gamma_{\rm red}$. For the GW signal, we adopt the domain wall annihilation scenario with four parameters in both datasets: the domain wall energy fraction $f_{w}(T_{\rm ann})$ at the annihilation temperature $T_{\rm ann}$, the domain wall annihilation temperature $T_{\rm ann}$ in units of GeV, the spectral index above the peak frequency $c_{1}$, and the parameter that represents the width of the spectral index $c_{2}$. The spectral $\gamma_{c}$ is fixed at 3 to respect the causality. In table 3, we provide the priors for the parameters of white and red noises, as well as the priors for parameters of the GW background signal from the annihilation of the domain walls.
@@
$E_{k}$ EFAC per backend/receiver system Uniform [0, $0,\leavevmode\nobreak\ 1$0] Single-pulsar analysis only
$Q_{k}$[s] EQUAD per backend/receiver system Log-uniform $8.5,\leavevmode\nobreak\ -5]$ Single-pulsar analysis only
$J_{k}$[s] ECORR per backend/receiver system Log-uniform $8.5,\leavevmode\nobreak\ -5]$ Single-pulsar analysis only
@@
$A_{\rm red}$ Red noise power-law amplitude Log-uniform $20,\leavevmode\nobreak\ -11]$ one parameter per pulsar
$\gamma_{\rm red}$ Red noise power-law spectral index Uniform $0,\leavevmode\nobreak\ 7]$ one parameter per pulsar
@@
$f_{w}(T_{\rm ann})$ Domain wall energy fraction at $T_{\rm ann}$ Log-Uniform [2,0][-2,0] one parameter per PTA
$T_{\rm ann}$ Domain wall annihilation temperature Log-Uniform $2.3,\leavevmode\nobreak\ 0]$ one parameter per PTA
$c_{1}$ GW spectral index above peak frequency Uniform $0.5,\leavevmode\nobreak\ 1.5]$ one parameter per PTA
$c_{2}$ Width of spectral index Uniform $0.5,\leavevmode\nobreak\ 2.5]$ one parameter per PTA
Table 4: Bayes estimators for the fit to NG15 and IPTA-DR2 datasets.
NG15
IPTA-DR2
${\rm Log}_{10}f_{w}(T_{\rm ann})$
$0.71\pm 0.17$
$0.93\pm 0.10$
${\rm Log}_{10}[T_{\rm ann}/\rm GeV]$
$0.70\pm 0.19$
$1.28\pm 0.15$
$c_{1}$
@@
[…]
@@
Appendix F Models for $N$ domain walls
To realize a model with $\mathcal{N}$ domains (and the domain wall number $N\sim\mathcal{N}$), we can consider the model with a $U(1)$ symmetry, which is softly broken by the term [21, 116]
$V(\Phi)\supset\frac{1}{M^{\mathcal{N}-4}}(\Phi^{\mathcal{N}}+\Phi^{\dagger\mathcal{N}})$ (F.1)
to a $Z_{\mathcal{N}}$ symmetry, where $\Phi$ is a complex scalar $\Phi$, $M$ has a mass dimension. For a renormalization model with $\mathcal{N}\leq 4$, the PBH production can be significant, and therefore, is strongly constrained.
One can consider $\mathcal{N}$ copies of complex scalars $\Phi_{i}$ (with $i=1,...,\mathcal{N}$). The global $U(1)^{\mathcal{N}}$ symmetry is spontaneously broken when each of $\Phi_{i}$ obtains a common vacuum expectation value. We can further explicitly break the $\mathcal{N}$ $U(1)$ symmetries to $\mathcal{N}-1$ discrete shift symmetries and one $U(1)$ symmetry by a term $\Phi_{i}^{\dagger}\Phi_{i+1}^{q}$, which then leads to the formation of $\mathcal{N}-1$ domains. Such symmetries have been naturally embedded in the framework of the clockwork axion models [117] and deconstruction of extradimensional grand unified theories (GUTs) [118, 119].
Another realization of the $Z_{\mathcal{N}}$ symmetry is the axion-like models (see, for example, Ref. [120]). Consider a hidden non-Abelian $SU(n)$ gauge sector consisting of a complex scalar $\Phi$ and a set of vector-like quarks. The potential of $\Phi$ possesses a global $U(1)$ symmetry, which is spontaneously broken by the Higgs mechanism. If the vector-like quarks are charged under the $U(1)$ transformation, then the quantum anomaly of $U(1)$ with respect to the $SU(n)$ gauge group can break the $U(1)$ symmetry to its $Z_{\mathcal{N}}$ subgroup via the instanton effect during the confinement of the $SU(n)$ symmetry, which gives rise to an axion-like periodic potential
$V(a)\simeq\Lambda_{n}^{4}\left(1-\cos\left(\frac{\mathcal{N}a}{v}\right)\right),$ (F.2)